OpenStudy (anonymous):

Solve for y 2+cos(3y)=x

OpenStudy (anonymous):

y= arccos(x-2)/3

OpenStudy (anonymous):

what?

geerky42 (geerky42):

2 + cos(3y) = x cos(3y) = x - 2 3y = \(\cos^{-1}(x - 2)\) \(\boxed{y = \dfrac{\cos^{-1}(x - 2)}{3}}\)

geerky42 (geerky42):

Is this clear?

OpenStudy (anonymous):

yes thank you

geerky42 (geerky42):

Glad we helped.